OpenStudy (anonymous):

Choose the correct simplification of f to the 9th power times h to the 23rd power all over f to the 3rd power times h to the 17th power. f12h6 1 over f to the 12th power times h to the 6th power f6h6 1 over f to the 6th power times h to the 6th power

OpenStudy (agent0smith):

\[\Large \frac{ f^9 h^{23} }{ f^3 h^{17}}\]

OpenStudy (anonymous):

what do i do after that?

OpenStudy (agent0smith):

Use this rule to simplify it\[\Large \frac{ x^a }{x^b } = x^{a-b}\]

OpenStudy (anonymous):

i got x^14-14

OpenStudy (agent0smith):

How'd you get that?

OpenStudy (anonymous):

sorry did it wrong hold on

OpenStudy (anonymous):

f6h6

OpenStudy (anonymous):

what about this one? Choose the correct simplification of 3 over x to the power of negative 7. x to the 7th power over 3 3 over x to the 7th power 3x7 Already simplified.

OpenStudy (agent0smith):

\[\Large \left( \frac{ 3 }{ x } \right)^{-7}\]

OpenStudy (anonymous):

so B?

OpenStudy (anonymous):

looks like its already simplified

OpenStudy (agent0smith):

It isn't, you can't have negative powers in simplified form.

OpenStudy (anonymous):

oh so its B?

OpenStudy (agent0smith):

Well you didn't say this wasn't the problem... \[\Large \left( \frac{ 3 }{ x } \right)^{-7}\] but it sounds more like \[\Large \frac{ 3 }{ x^{-7} }\]

OpenStudy (anonymous):

it is \[\frac{ 3 }{ x^-7 }\]

OpenStudy (agent0smith):

For negative exponents \[\Large x^{-n} = \frac{ 1 }{ x^n }\] and also \[\Large \frac{ 1 }{ x^{-n} } = x^n\]

OpenStudy (anonymous):

its 3x^7

OpenStudy (agent0smith):

Yep :)

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

I have one more and then I'm done, Which expression is equivalent to the area of metal sheet required to make this square-shaped traffic sign? A square shaped traffic sign is shown with length of one side labeled as x + 1. x^2 + 2x + 1 x^2 + x + 1 x^2 + 2x x^2 + 1

OpenStudy (anonymous):

I think it's A but I'm not sure

OpenStudy (agent0smith):

If the side length is x+1, then the area is the (side length)^2, or (x+1)^2. (x+1)^2 = (x+1)(x+1) =